Multiply the radical expressions.
a.) √3 × √15
b.) √8 × √12

Found 3 solutions by ReadingBoosters, kevwill, stanbon:
Answer by ReadingBoosters(3246)

duplicate
.....................
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.

Answer by kevwill(135)
Answer by stanbon(75887)

Multiply the radical expressions.
a.) √3 × √15
-----
= sqrt(45) = sqrt(9*5) = sqrt(9)*sqrt(5) = 3sqrt(5)
--------------------------
b.) √8 × √12 = 2sqrt(2) * 2sqrt(3) = 4sqrt(6)
===============================================
Cheers,
Stan H.